# QVGen: Pushing the Limit of Quantized Video Generative Models

### v4

@@ 4.4 Efficiency Discussion
We report the per-step latency of W4A4 DiT components on one $A800$ GPU in Fig. 6 (b). Adapted from the CUDA kernel implementation by Ashkboos et al. (2024), W4A4 QVGen achieves $1.21\times$ and $1.44\times$ speedups for Wan 1.31.3B and 1414B, respectively. Besides, it exhibits $\sim 4\times$ memory savings compared to the BF16 format, as shown in Fig. 6 (a). Nevertheless, we believe that the acceleration ratio could be further improved with advanced kernel-fusion techniques and optimization for specific tensor shapes in these models, which we leave to future work. In addition, it is worth noting that our QVGen adheres to standard uniform quantization, enabling drop-in deployment with existing W4A4 kernels across various devices.
@@ 5 Related Work
Model quantization. Quantization (Jacob et al., 2017) is a predominant technique for minimizing storage and accelerating inference. It can be categorized into post-training quantization (PTQ) (Nagel et al., 2020) and quantization-aware training (QAT) (van Baalen et al., 2020). PTQ compresses models without re-training, making it fast and data-efficient. Nevertheless, it may result in suboptimal performance, especially under ultra-low bit-width (e.g., 33/44-bit). Conversely, QAT applies quantization during training or finetuning and typically achieves higher compression rates with less performance degradation. For DMs, previous quantization research (Li et al., 2023a; He et al., 2023; Huang et al., 2024a; 2025; So et al., 2023; Wu et al., 2024a; Shang et al., 2023; Wang et al., 2024) primarily focuses on image generation. Video DMs, which incorporate complex temporal and spatial modeling, are still challenging for low‑bit quantization. QVD (Tian et al., 2024) and Q-DiT (Chen et al., 2024b) first apply PTQ to convolution-based (Xu et al., 2023b; Guo et al., 2024) and DiT-based (Zheng et al., 2024) video DMs, respectively. Furthermore, ViDiT‑Q (Zhao et al., 2025a) employs mixed‑precision and fine‑grained PTQ to improve performance. However, it experiences video quality loss with 8‑bit activation quantization and has not been extended to more advanced models (Yang et al., 2025; WanTeam et al., 2025). Consequently, QAT for advanced video DMs is urgently needed.
In this work, we identify and address the ineffective low-bit QAT for the video DM. Our proposed method significantly enhances model performance and incurs zero inference overhead in 33/44-bit settings. Moreover, our framework is orthogonal to existing QAT methods, which target gradient estimation (Gong et al., 2019), oscillation reduction (Nagel et al., 2022), etc. We provide more related work about diffusion models in Sec. A.
@@ 6 Conclusions and Limitations
Appendix A More Related Work
Video diffusion models. Building upon the remarkable success of diffusion models (DMs) (Ho et al., 2020a; Song et al., 2021a; Chen et al., 2024a) in image generation (Labs, 2024; Xie et al., 2025), the exploration in the field of video generation (Yang et al., 2025; WanTeam et al., 2025; Kong et al., 2025) is also becoming popular. In contrast to convolution‑based diffusion models (Ho et al., 2022; Blattmann et al., 2023), the success of OpenAI Sora (OpenAI, 2024) has spurred researchers to adopt the diffusion transformer (DiT) (Peebles & Xie, 2023) architecture and scale it up for high‑quality video generation. However, advanced video DiTs (WanTeam et al., 2025; Yang et al., 2025; Kong et al., 2025) often involve billions of parameters, lengthy multi‑step denoising, and intensive computation over long frame sequences. This results in substantial time and memory overhead, which limits their practical deployment. To enable faster video generation, some works have introduced step‑distillation (Yin et al., 2025; Lin et al., 2025) on pre-trained models to shorten the denoising trajectory. Others focus on efficient attention (Zhang et al., 2025a; Xi et al., 2025), feature caching (Lv et al., 2025; Zou et al., 2025), or parallel inference (Fang et al., 2024b; a) to accelerate per-step computations. Moreover, to achieve memory-efficient inference, existing research has explored efficient architecture design (Wu et al., 2024b; Liu et al., 2025), structure pruning (Ben Yahia et al., 2024; Zhao et al., 2025b), and model quantization (Zhao et al., 2025a; Chen et al., 2024b; Tian et al., 2024). These methods aim to achieve both model size and computational cost reduction.
Appendix B Alogrithm of QVGen
@@
Appendix C Proof of Thm. 3.1
Assumption C.1.
@@
Considering Ass. C.2, we can further relax the above inequality to:
@@ Appendix D More Experimental Details
Baselines. For QAT baselines, we employ the same settings as our QVGen, e.g., training iterations, batch size, and optimizer, to make a fair comparison. For PTQ baselines, we adopt bit-width in $\{2,\ldots,8\}$ for the mixed-precision strategy proposed in ViDiT-Q (Zhao et al., 2025a). For W4A4 group-wise SVDQuant (Li et al., 2025), we retain $16$-bit precision for linear layers involved in adaptive normalization, embedding layers, and the key and value projections in cross-attention. For both W4A4 and W4A6 SVDQuant, we apply SmoothQuant (Xiao et al., 2024) as a pre-processing step, and GPTQ (Frantar et al., 2023) as a post-processing step. Except as noted above, we only quantize all linear layers for a given video DM. For the attention module, we adopt full-precision flash-attention (Dao et al., 2022) to speed up inference, which is a common practice. It is also worth noting that, since we only quantize linear layers and employ flash-attention for the attention modules, the “Naive” baseline (see Tab. O) is equivalent to Q-DM (Li et al., 2023b) in this paper.
Training. Before QAT, we resize and center-crop the input frames to match the evaluation resolution, except for Wan 1414B, which uses $480p$ during QAT to reduce training costs. We then sample video clips containing $49$ frames at equal-frame intervals. During QAT, Wan 1414(WanTeam et al., 2025) and CogVideoX1.5-55(Yang et al., 2025) are trained using PyTorch FSDP (Zhao et al., 2023b), while other diffusion models are trained with DeepSpeed ZeRO-2 (Rajbhandari et al., 2020). Specifically, we adopt a warm-up phase spanning $\frac{1}{10}$ of the total training epochs, and set the global batch size to $48$ for Wan 1.31.3B and $64$ for all other models. A cosine annealing schedule is applied to the learning rate  [^6] This mentioned learning rate is applied to both model weights and the introduced $\Phi$., initialized at $3\times 10^{-5}$ for moderate-sized models ($\leq$22B), $5\times 10^{-5}$ for CogVideoX1.5-55B, and $10^{-5}$ for Wan 1414B. For weight quantization parameters, we adopt LSQ (Esser et al., 2020b) with an initial learning rate of $3\times 10^{-5}$. As described in the preliminary section of the main text, we use the straight-through estimator (STE) (Bengio et al., 2013) to ensure the differentiability of the quantization process. To be noted, when the remaining rank $r<\frac{1}{\lambda}$, we directly apply a cosine annealing function (i.e., ${\bm{\gamma}}=[u]_{n\times r}$) to gradually shrink the remaining ${\mathbf{W}}_{\Phi}$ to $\varnothing$. Moreover, the training days are summarized in Tab. G.
Table G: #GPU days ($H100$) across different models. We present the increased time of QVGen compared with the naive QAT in a KD-based manner in red subscripts. QVGen, which uses $r=32$ in this paper, incurs negligible time overhead but significantly higher performance (see Tab. O) than the naive method.
@@
Evaluation. For VBench (Huang et al., 2024b), We test $1\sim 2$B models on $8\times H100$ GPUs. For huge DMs, we evaluate CogVideoX1.5-55B on $64\times H100$ GPUs and Wan 1414B on $128\times H100$ GPUs for both VBench and VBench-2.0 (Zheng et al., 2025a). The batch size is set to one per GPU, and each run completes within one day. Besides, we sample $5$ videos for each unaugmented text prompt across the $8$ dimensions in VBench. For VBench-2.0 (Zheng et al., 2025a), we generate $3$ videos per augmented text prompt across all dimensions, except for the Diversity dimension, where we generate $20$ videos for each prompt.
@@ Appendix E Comparison with Baselines under relatively higher bit-width
Table H: Performance comparison across different methods on VBench under W6A6 quantization for Wan 1.31.3B.
Method
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Besides 33/44-bit settings, we also conduct experiments under W6A6. As shown in Tab. H, our QVGen achieves full-precision comparable or even better performance than BF16 models. Since these settings are not challenging enough for QAT baselines, which also demonstrate satisfactory performance, our QVGen only shows moderate improvements.

### v5

@@ 4.4 Efficiency Discussion
Inference efficiency. We report the per-step latency of W4A4 DiT components on one $A800$ GPU in Fig. 6 (b). Adapted from the CUDA kernel implementation by Ashkboos et al. (2024), W4A4 QVGen achieves $1.21\times$ and $1.44\times$ speedups for Wan 1.31.3B and 1414B, respectively.
@@
Besides, it exhibits ${\sim}4\times$ memory savings compared to the BF16 format, as shown in Fig. 6 (a). Nevertheless, we believe that the acceleration ratio could be further improved with advanced kernel-fusion techniques and optimization for specific tensor shapes in these models, which we leave to future work (More discussion can be found in Secs. N-O). In addition, it is worth noting that our QVGen adheres to standard uniform quantization, enabling drop-in deployment with existing W4A4 kernels across various devices.
@@
Table 7: Training costs across different methods and models.
[…]
LSQ Q-DM EfficientDM QVGen LSQ Q-DM EfficientDM QVGen
CogVideoX-22B 8.64 9.30 8.97 9.44 62.78 67.26 44.27 67.93
Wan 1.31.3B 9.92 10.92 10.68 11.11 63.04 66.15 42.74 66.67
Training efficiency. Moreover, we conduct a comparative training efficiency analysis across different models and QAT baselines, as demonstrated in Tab. 7. LSQ (Esser et al., 2020b) does not use distillation-based QAT (i.e., no teacher forward pass) and is therefore the fastest to train. EfficientDM (He et al., 2024) updates only LoRA parameters, which substantially reduces optimizer-state memory on GPUs. We believe such a strategy in EfficientDM can be combined with our method to further improve training efficiency, and we plan to explore this in future work. Relative to distillation-based QAT (i.e., Q-DM (Li et al., 2023b)), our method with low-rank $\Phi$ adds only ${\sim}1.02\times$ GPU-days and ${\sim}1.01\times$ peak GPU memory for Wan 1.31.3B model. To be noted, all baselines greatly fall short of our method in final performance, as illustrated in Sec. 4.2. In future work, we will delve into improving QVGen’s training efficiency while maintaining its strong performance.
@@ 5 Related Work
Video diffusion models. Building upon the remarkable success of diffusion models (DMs) (Ho et al., 2020a; Song et al., 2021a; Chen et al., 2024a) in image generation (Labs, 2024; Xie et al., 2025), the exploration in the field of video generation (Yang et al., 2025; WanTeam et al., 2025; Kong et al., 2025) is also becoming popular. In contrast to convolution‑based diffusion models (Ho et al., 2022; Blattmann et al., 2023), the success of OpenAI Sora (OpenAI, 2024) has spurred researchers to adopt the diffusion transformer (DiT) (Peebles & Xie, 2023) architecture and scale it up for high‑quality video generation. However, advanced video DiTs (WanTeam et al., 2025; Yang et al., 2025; Kong et al., 2025) often involve billions of parameters, lengthy multi‑step denoising, and intensive computation over long frame sequences. This results in substantial time and memory overhead, which limits their practical deployment. To enable faster video generation, some works have introduced step‑distillation (Yin et al., 2025; Lin et al., 2025) on pre-trained models to shorten the denoising trajectory. Others focus on efficient attention (Zhang et al., 2025a; Xi et al., 2025; Huang et al., 2025a), feature caching (Lv et al., 2025; Zou et al., 2025; Huang et al., 2024b), or parallel inference (Fang et al., 2024c; b) to accelerate per-step computations. Moreover, to achieve memory-efficient inference, existing research has explored efficient architecture design (Wu et al., 2024b; Liu et al., 2025), structure pruning (Ben Yahia et al., 2024; Zhao et al., 2025b), and model quantization (Zhao et al., 2025a; Chen et al., 2024b; Tian et al., 2024). These methods aim to achieve both model size and computational cost reduction.
Model quantization. Quantization (Jacob et al., 2017) is a predominant technique for minimizing storage and accelerating inference. It can be categorized into post-training quantization (PTQ) (Nagel et al., 2020) and quantization-aware training (QAT) (van Baalen et al., 2020). PTQ compresses models without re-training, making it fast and data-efficient. Nevertheless, it may result in suboptimal performance, especially under ultra-low bit-width (e.g., 33/44-bit). Conversely, QAT applies quantization during training or finetuning and typically achieves higher compression rates with less performance degradation. For DMs, previous quantization research (Li et al., 2023a; He et al., 2023; Huang et al., 2024a; 2025b; So et al., 2023; Wu et al., 2024a; Shang et al., 2023; Wang et al., 2024) primarily focuses on image generation. Video DMs, which incorporate complex temporal and spatial modeling, are still challenging for low‑bit quantization. QVD (Tian et al., 2024) and Q-DiT (Chen et al., 2024b) first apply PTQ to convolution-based (Xu et al., 2023b; Guo et al., 2024) and DiT-based (Zheng et al., 2024) video DMs, respectively. Furthermore, ViDiT‑Q (Zhao et al., 2025a) employs mixed‑precision and fine‑grained PTQ to improve performance. However, it experiences video quality loss with 8‑bit activation quantization and has not been extended to more advanced models (Yang et al., 2025; WanTeam et al., 2025). Consequently, QAT for advanced video DMs is urgently needed.
In this work, we identify and address the ineffective low-bit QAT for the video DM. Our proposed method significantly enhances model performance and incurs zero inference overhead in 33/44-bit settings. Moreover, our framework is orthogonal to existing QAT methods, which target gradient estimation (Gong et al., 2019), oscillation reduction (Nagel et al., 2022), etc.
@@ 6 Conclusions and Limitations
Acknowledgement
This work was supported by the National Natural Science Foundation of China (Nos. 62476018), and the Postdoctoral Fellowship Program of CPSF (No. BX20250487). This work was also supported by the Hong Kong Research Grants Council under the Areas of Excellence scheme grant AoE/E-601/22-R and NSFC/RGC Collaborative Research Scheme grant CRS_HKUST603/22.
@@ Acknowledgement
Appendix A Alogrithm of QVGen
@@
Appendix B Proof of Thm. 3.1
Assumption B.1.
@@
Considering Assm. B.2, we can further relax the above inequality to:
@@
Appendix C Convergence Analysis without Requiring Convexity
For completeness, we also provide a nonconvex convergence result for video DMs under standard smoothness assumptions. Here we consider a (possibly nonconvex) training objective $F:\mathbb{S}^{d}\to\mathbb{R}$ and the gradient descent updates
$\begin{array}[]{ll}{\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}{\bm{g}}_{t},\qquad{\bm{g}}_{t}:=\nabla F({\bm{\theta}}_{t}),\end{array}$ (I)
where $t$ indexes the optimization iterations. Regret analysis in Sec. B follows the standard online learning setting, where the per-step loss $f_{t}$ varies with $t$. In contrast, nonconvex convergence analysis (Bottou et al., 2018; Ghadimi & Lan, 2013) uses a fixed deep-learning objective $F$ across iterations, so no subscript is needed, and convergence is certified by a vanishing minimum gradient norm.
Assumption C.1 (Smoothness).
The function $F$ is $L$-smooth, i.e., for all ${\bm{\theta}},{\bm{\theta}}^{\prime}\in\mathbb{S}^{d}$,
$\begin{array}[]{ll}\|\nabla F({\bm{\theta}})-\nabla F({\bm{\theta}}^{\prime})\|_{2}\leq L\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|_{2}.\end{array}$ (J)
Equivalently, $F$ satisfies the standard descent inequality:
$\begin{array}[]{ll}F({\bm{\theta}}^{\prime})\leq F({\bm{\theta}})+\nabla F({\bm{\theta}})^{\top}({\bm{\theta}}^{\prime}-{\bm{\theta}})+\frac{L}{2}\|{\bm{\theta}}^{\prime}-{\bm{\theta}}\|_{2}^{2}.\end{array}$ (K)
Assumption C.2 (Learning rate bounds).
The learning rates satisfy
$\begin{array}[]{ll}0<\eta^{m}\leq\eta_{t}\leq\eta^{M}\leq\frac{1}{L}\quad\text{for all }t,\end{array}$ (L)
where $\eta^{m}$ and $\eta^{M}$ are the minimum and maximum learning rates [^6] We employ the same learning rate for the entire quantized DM for simplicity. across iterations, respectively.
Assumption C.3 (Lower-bounded objective).
There exists $F^{\star}$ such that $F({\bm{\theta}})\geq F^{\star}$ for all ${\bm{\theta}}\in\mathbb{S}^{d}$.
Theorem C.4 (Convergence to a first-order stationary point).
Under Assms. C.1C.3, the iterates of Eq. (I) satisfy
$\begin{array}[]{ll}\frac{1}{T}\sum_{t=1}^{T}\|{\bm{g}}_{t}\|_{2}^{2}\;\leq\;\frac{2\big(F({\bm{\theta}}_{1})-F^{\star}\big)}{T\,\eta^{m}}.\end{array}$ (M)
Consequently,
$\begin{array}[]{ll}\min_{1\leq t\leq T}\|{\bm{g}}_{t}\|_{2}^{2}\;\leq\;\frac{1}{T}\sum_{t=1}^{T}\|{\bm{g}}_{t}\|_{2}^{2}\xrightarrow{T\to\infty}0.\end{array}$ (N)
Thus, the iterates converge in the standard nonconvex sense to first-order stationary points, in the sense that there exists an iterate with arbitrarily small gradient norm when $T$ is sufficiently large.
Proof.
By $L$-smoothness (the descent inequality) and the update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}{\bm{g}}_{t}$,
$\begin{array}[]{ll}F({\bm{\theta}}_{t+1})&\leq F({\bm{\theta}}_{t})+\nabla F({\bm{\theta}}_{t})^{\top}({\bm{\theta}}_{t+1}-{\bm{\theta}}_{t})+\frac{L}{2}\|{\bm{\theta}}_{t+1}-{\bm{\theta}}_{t}\|_{2}^{2}\\ &=F({\bm{\theta}}_{t})-\eta_{t}\|{\bm{g}}_{t}\|_{2}^{2}+\frac{L}{2}\eta_{t}^{2}\|{\bm{g}}_{t}\|_{2}^{2}=F({\bm{\theta}}_{t})-\eta_{t}\Big(1-\frac{L\eta_{t}}{2}\Big)\|{\bm{g}}_{t}\|_{2}^{2}.\end{array}$ (O)
Since $\eta_{t}\leq 1/L$, we obtain $1-L\eta_{t}/2\geq 1/2$, hence
$\begin{array}[]{ll}F({\bm{\theta}}_{t+1})\leq F({\bm{\theta}}_{t})-\frac{\eta_{t}}{2}\|{\bm{g}}_{t}\|_{2}^{2}.\end{array}$ (P)
Summing over $t\in\{1,\ldots,T\}$ and using $F({\bm{\theta}}_{T+1})\geq F^{\star}$ yields
$\begin{array}[]{ll}\sum_{t=1}^{T}\frac{\eta_{t}}{2}\|{\bm{g}}_{t}\|_{2}^{2}\leq F({\bm{\theta}}_{1})-F^{\star}.\end{array}$ (Q)
With $\eta_{t}\geq\eta^{m}$, we obtain
$\begin{array}[]{ll}\frac{1}{T}\sum_{t=1}^{T}\|{\bm{g}}_{t}\|_{2}^{2}\leq\frac{2\big(F({\bm{\theta}}_{1})-F^{\star}\big)}{T\,\eta^{m}},\end{array}$ (R)
which proves Eq. (M). Moreover, we have
$\begin{array}[]{ll}\min_{1\leq t\leq T}\|{\bm{g}}_{t}\|_{2}^{2}\;\leq\;\frac{1}{T}\sum_{t=1}^{T}\|{\bm{g}}_{t}\|_{2}^{2}\;\leq\;\frac{2\big(F({\bm{\theta}}_{1})-F^{\star}\big)}{T\,\eta^{m}}\xrightarrow{T\to\infty}0.\end{array}$ (S)
Remark C.5 (Connection to our convex analysis).
Thm. C.4 shows that, for any finite $T$, the convergence behavior (i.e., Eq. (N)) of QAT is determined by the average gradient norm $\frac{1}{T}\sum_{t=1}^{T}\|{\bm{g}}_{t}\|_{2}^{2}$. Since this quantity admits an $\mathcal{O}(1/T)$ upper bound under smoothness and bounded learning rates, reducing $\|{\bm{g}}_{t}\|_{2}$ during training directly tightens the bound and improves the finite-step convergence of QAT. This matches our convex regret analysis, where the same average gradient norm appears as the core term controlling convergence. Thus, reducing the gradient norm is essential for improving QAT optimization in both settings.
@@ Appendix D More Experimental Details
Baselines. For QAT baselines, we employ the same settings as our QVGen, e.g., training iterations, batch size, and optimizer, to make a fair comparison. For PTQ baselines, we adopt bit-width in $\{2,\ldots,8\}$ for the mixed-precision strategy proposed in ViDiT-Q (Zhao et al., 2025a). For W4A4 group-wise SVDQuant (Li et al., 2025), we retain $16$-bit precision for linear layers involved in adaptive normalization, embedding layers, and the key and value projections in cross-attention. For both W4A4 and W4A6 SVDQuant, we apply SmoothQuant (Xiao et al., 2024) as a pre-processing step, and GPTQ (Frantar et al., 2023) as a post-processing step. Except as noted above, we only quantize all linear layers for a given video DM. For the attention module, we adopt full-precision flash-attention (Dao et al., 2022) to speed up inference, which is a common practice. It is also worth noting that, since we only quantize linear layers and employ flash-attention for the attention modules, the “Naive” baseline (see Tab. K) is equivalent to Q-DM (Li et al., 2023b) in this paper.
Training. Before QAT, we resize and center-crop the input frames to match the evaluation resolution, except for Wan 1414B, which uses $480p$ during QAT to reduce training costs. We then sample video clips containing $49$ frames at equal-frame intervals. During QAT, Wan 1414(WanTeam et al., 2025) and CogVideoX1.5-55(Yang et al., 2025) are trained using PyTorch FSDP (Zhao et al., 2023b), while other diffusion models are trained with DeepSpeed ZeRO-2 (Rajbhandari et al., 2020). Specifically, we adopt a warm-up phase spanning $\frac{1}{10}$ of the total training epochs, and set the global batch size to $48$ for Wan 1.31.3B and $64$ for all other models. A cosine annealing schedule is applied to the learning rate  [^7] This mentioned learning rate is applied to both model weights and the introduced $\Phi$., initialized at $3\times 10^{-5}$ for moderate-sized models ($\leq$22B), $5\times 10^{-5}$ for CogVideoX1.5-55B, and $10^{-5}$ for Wan 1414B. For weight quantization parameters, we adopt LSQ (Esser et al., 2020b) with an initial learning rate of $3\times 10^{-5}$. As described in the preliminary section of the main text, we use the straight-through estimator (STE) (Bengio et al., 2013) to ensure the differentiability of the quantization process. To be noted, when the remaining rank $r<\frac{1}{\lambda}$, we directly apply a cosine annealing function (i.e., ${\bm{\gamma}}=[u]_{n\times r}$) to gradually shrink the remaining ${\mathbf{W}}_{\Phi}$ to $\varnothing$. Moreover, the training days are summarized in Tab. A.
Table A: #GPU days ($H100$) across different models. We present the increased time of QVGen compared with the naive QAT in a KD-based manner in red subscripts. QVGen, which uses $r=32$ in this paper, incurs negligible time overhead but significantly higher performance (see Tab. K) than the naive method.
@@
Evaluation. For VBench (Huang et al., 2024c), we test $1{\sim}2$B models on $8{\times}H100$ GPUs. For huge DMs, we evaluate CogVideoX1.5-55B on $64{\times}H100$ GPUs and Wan 1414B on $128{\times}H100$ GPUs for both VBench and VBench-2.0 (Zheng et al., 2025a). The batch size is set to one per GPU, and each run completes within one day. Besides, we sample $5$ videos for each unaugmented text prompt across the $8$ dimensions in VBench. For VBench-2.0 (Zheng et al., 2025a), we generate $3$ videos per augmented text prompt across all dimensions, except for the Diversity dimension, where we generate $20$ videos for each prompt.
@@ Appendix E Comparison with Baselines under relatively higher bit-width
Table B: Performance comparison across different methods on VBench under W6A6 quantization for Wan 1.31.3B.
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
Besides 33/44-bit settings, we also conduct experiments under W6A6. As shown in Tab. B, our QVGen achieves full-precision comparable or even better performance than BF16 models. Since these settings are not challenging enough for QAT baselines, which also demonstrate satisfactory performance, our QVGen only shows moderate improvements.